OpenStudy (anonymous):

Help!!!!! Determine the delivery radius for your shop. Draw a point on a coordinate plane where your shop will be located. Create two different radii lengths from your shop, and construct the circles that represent each delivery area. How much area will each delivery radius cover? Write the equation for each circle created. Please show your work for all calculations.

OpenStudy (anonymous):

I have to create a pizza shop. This is what I have so far. I'm very confused.

OpenStudy (anonymous):

Is your first radius 4.5 units long?

OpenStudy (anonymous):

Yes, it is.

OpenStudy (anonymous):

Ok, so you need to create another circle with a different radius length and then solve for the areas of the two different circles. The formula for area of a circle is pi(r)^2 where r represents the length of the radius. So, your first equation would be Pi(4.5)^2

OpenStudy (anonymous):

\[3.14(4.5)^2=\]

OpenStudy (anonymous):

The equation of a circle in x-y coordinates is: (x-g)² + (y-h)² = R² where the radius of the circle is R and the center of the circle is at (x,y) = (g, h) (g,h) would be the coordinates of the location of you shop, and R would be the radius of the delivery area. Assuming that your shop is located at the center of the coordinate system, (x,y) = (0,0), and your delivery radius is 10 miles, the equation of the circle representing that delivery area would be: (x-0)² + (y-0)² = 10² or x² + y² = 100 where x and y are measured in miles.

OpenStudy (anonymous):

I'm still struggling with the equation creator lol

OpenStudy (anonymous):

@ayeeraee is there is any doubts

OpenStudy (anonymous):

@best.shakir not yet, I'm gonna try this out

OpenStudy (anonymous):

he's right

OpenStudy (anonymous):

10.125*2

OpenStudy (anonymous):

@chrisisbad92 how should I construct the new circles?

OpenStudy (anonymous):

However you like. Decide a new place for the midpoint of your 2nd circle and make sure the radius is either longer or shorter than 4.5.

OpenStudy (anonymous):

I'd assume you don't want them to intersect either, so using a different quadrant woud probably be best.

OpenStudy (anonymous):

Okay, I'm gonna try to work out the equation. Hold on..

OpenStudy (anonymous):

hold on...shouldn't the 2 circles be concentric? they didn't mention to change the location of ur shop. so, for the second circle, keep the centre same and only change the value of ur radius.

OpenStudy (anonymous):

You're right Vaidehi09. The problem doesn't state two different locations for the midpoint.

OpenStudy (anonymous):

try 10.125*2 this may help chrisisbad92

OpenStudy (anonymous):

@Vaidehi09, okay, so I can two radii pointing in different directions?

OpenStudy (anonymous):

i don't really get what u mean by that...but this is how ur figure should look like: